OpenStudy (anonymous):

Solve approx to 5th decimal place. 7^(2x+1) = 3^(x+2) x= ?

OpenStudy (anonymous):

Distribute 7 and 3 out. 14x+7 = 3x+6 Put the variable on one side of the equal sign and the constants on the other. Ex: 11x = -1 x = -1/11 Use a calculator and round to appropriate decimal places.

OpenStudy (anonymous):

this is incorrect the '^' indicates an exponent - not multiply to solve this take logs ln 7^(2x+1) = ln 3^(x+2) (2x+1) ln 7 = (x+2) ln 3 (2x+1) / (x+2) = ln 3 / ln 7 = 0.564575 2x - 0.564575x = 2 * 0.564575 - 1 1.435425x = 0.12915 x = 0.08997 to 5 dec.places

OpenStudy (anonymous):

Oops, sorry! I didn't read it carefully :(